# Disentangling the Potential Impacts of Papers into Diffusion, Conformity, and Contribution Values

### v3

@@ 3. Methodology
In this section, we present a detailed exposition of our proposed model, DPPDCC. Initially, we formulate the constructed dynamic heterogeneous and the potential impact prediction task. Following this, we provide an in-depth elucidation of DPPDCC’s core components: the Citation-aware GNN Encoder (CGE) and the Popularity-aware Disentanglement Module (PDM), as depicted in Figure 2. The CGE integrates multiple layers of structural and temporal encoders, aimed at capturing intricate high-order temporal and structural information within the network. Concurrently, the PDM comprises three distinct modules, each targeting a distinct perspective of citation increment, all trained in parallel. Lastly, we delineate the training and optimization strategies employed for DPPDCC.
Figure 2. The overall architecture of DPPDCC. Initially, DPPDCC processes the dynamic heterogeneous graph of the target paper using the Citation-aware GNN Encoder. Within this encoder, the structural component utilizes CompGAT to extract comparative and co-cited/citing information between papers at each time step. This information is then aggregated evolutionarily across snapshots via type-specific attention readout. Subsequently, aggregated snapshot information from different time steps is interacted in the temporal encoder. By stacking multiple layers of the structural encoder and temporal encoder, it enables CGE to encode complex high-order temporal and structural information effectively. Subsequent to obtaining the target paper’s representation, DPPDCC proceeds to disentangle its citation increment into diffusion, conformity, and contribution values through three corresponding auxiliary tasks. The diffusion module augments the graphs by employing weighted node dropping, facilitating the identification of the diffusion process by contrasting augmented views with the original graph. The conformity module quantifies conformity levels by binning accumulated citations. Finally, the contribution module introduces orthogonal constraints, prompting each module to learn distinct information from diverse perspectives.
@@ 3.1. Problem Formulation and Notations
Before presenting our approach, we first introduce the problem formulation and notations. In practice, the citation network comprises diverse entities and relations that evolve dynamically each year. To capture the entity-aware and time-variant characteristics, of each target paper, we construct a dynamic heterogeneous graph from the citation network to predict its potential impact.
@@ 3.1.1. Dynamic Heterogeneous Graph
Given a target paper at the time step $t$, its heterogeneous subgraph is defined as $G^{t}=(\nu^{t},\epsilon^{t},\phi,\varphi)$. Here, $\nu^{t}$ denotes the set of nodes surrounding the target paper at the time step $t$. And $\epsilon^{t}$ represents the set of edges among $\nu^{t}$. Each node $\nu$ and edge $\epsilon$ are associated with their type mapping functions $\phi:\nu\rightarrow U$ and $\varphi:\epsilon\rightarrow R$, where $U$ and $R$ denote the types of nodes and edges (see Figure 1 (c)), respectively.
The backbone of the heterogeneous subgraph primarily consists of the cited and citing relations. We collect $k$-hop cited/citing neighbors of the target paper at the specific time $t$. For each $i$-th hop, to maintain the most recent and significant information, we select the $K_{i}$ newest published and then most highly cited papers, for references and citations respectively. To address the potential lack of global information, we incorporate metadata nodes associated with the papers (authors, venue, and publication time). Prior to the observation time point, we consider a total of $T$ such subgraphs, each serving as a snapshot captured at distinct time steps. These subgraphs organized into a sequence collectively constitute the dynamic heterogeneous graph $G$ of the target paper.
@@ 3.1.2. Potential Impact Prediction
We assess the potential impact of the given target paper by estimating its citation increment. It is quantified as the increased citation number $\Delta$ years after the observation time point. Given a paper, we try to learn a function $f(\cdot)$ to predict the number of citation increments.
@@ 3.2. Initialization
The initial embedding of the paper node is derived from the text embedding obtained by encoding the combination of title and abstract. Metadata node embeddings in each snapshot are the average embeddings over all papers associated with them in the global citation network at the corresponding time step. Specifically, the extraction of text embeddings is accomplished using Sentence-BERT (Reimers and Gurevych, 2019). For $i$-th node at the $t$-th time step, we have:
(1) $\begin{split}&\bm{h}_{ti,paper}^{(0)}=TextEncoder([w_{1},w_{2},...,w_{n^{w}}])\\ &\bm{h}_{ti,u}^{(0)}=avgpool([\bm{h}_{t1,paper}^{(0)},\bm{h}_{t2,paper}^{(0)},...,\bm{h}_{tm_{p},paper}^{(0)}]),\\ &u\in\{author,venue,time\}\end{split}$
where $u$ denotes the node type, $TextEncoder$ denotes the text representation model, and $m_{p}$ is the count of metadata node’s connected papers.
To initialize the 00-th layer representation of the snapshot of the $i$-th target paper at the $t$-th time step $\bm{h}_{ti,s}^{(0)}$, we mainly consider the contextual information from the corresponding time embedding $\bm{h}_{t,time}^{(0)}$ at time step $t$. Since the time embedding only preserves the global features, we utilize a gate module to approach specific information based on the $i$-th target paper embedding $\bm{h}_{i,p}^{(0)}$. The snapshot representation $\bm{h}_{ti,s}^{(l)}$ of $l$-th layer will then be the proxy to aggregate the particular heterogeneous subgraph:
(2) $\begin{split}&\bm{h}_{ti,s}^{(0)}=\bm{a}_{ti}\times\bm{h}_{t,time}^{(0)}\\ &\bm{a}_{ti}=\mathrm{sigmoid}(\bm{W}^{ipt}\bm{h}_{t,time}^{(0)}+\bm{W}^{tgt}\bm{h}_{i,p}^{(0)})\end{split}$
@@ 3.3. Citation-aware GNN Encoder
Citation-aware GNN Encoder (CGE) aims to capture distinct structural and temporal features from the dynamic heterogeneous graph, which is facilitated by leveraging knowledge diffusion theories derived from bibliometrics. To be specific, the backbone of CGE consists of modified R-GCNs as the structural encoder and a Transformer Encoder as the temporal encoder. We stack $L$ integrated layers to encode them alternately instead of separate modeling. This allows us to exploit higher-order interactions between space and time. To elaborate, we first employ distinct Citation-aware R-GCNs to model structural information for subgraphs at different time points individually. Afterwards, a Type-specific Attention Snapshot Readout is employed to summarize the information from these subgraphs established on the snapshot representations. Simultaneously, the snapshot information mixer integrates global and temporal information from snapshot representations into the local and structural representations of paper nodes, fostering interactions among various feature spaces. The aggregated snapshot representations are then input into the temporal encoder (Encoder of the Transformer) to capture the intricate temporal relationships between different time points. By integrating structural and temporal encoding, our approach could extract valuable insights from the dynamic heterogeneous graph.
@@ 3.3.1. Citation-aware R-GCNs
Citation-aware R-GCNs, as the core component of the structural encoder, are utilized to encode node hidden states $\bm{h}$ within the dynamic heterogeneous graph. Following the manner of R-GCN (Schlichtkrull et al., 2018), we deliver message passing in each relation separately for modeling attentively. Moreover, we replace the original GCNs in R-GCN with more powerful GNN modules to sensitively encode intricate information.
@@
where $\mathcal{R}$ is the set of all edges in the graph, $g_{r}$ is the subgraph with relation $r$, and $AGG$ is the relation aggregation function. Here, we adopt sum-pooling as the relation aggregation function for efficiency and performance.
For ”cites” and ”is cited by” edge subgraphs of paper nodes, inspired by Disruption-index (Wu et al., 2019), we introduce co-cited/citing strengths to integrate multi-hop citation information. The Disruption-index categorizes papers as disruptive or developmental based on the metric related to their betweenness centrality. Disruptive papers play a critical role in altering typical citation diffusion. We adapt this subgraph-focused metric to estimate node-pair closeness by considering co-cited/citing strengths. We aim to discern the varying significance of citation behaviors during message passing and distinguish crucial references and citations within ”is cited by” and ”cites” edges for the target paper. On one side, within ”is cited by” edges, when the source (reference) and target (citation) papers share more similar references, the influence of the source paper on the target paper might diminish. From the viewpoint of the target paper, the source paper could be considered as merely developmental, exerting a potentially negative impact on message passing. Moreover, since the information from this source paper is already conveyed through other original disruptive references, the negative co-citing effect might amplify attention toward these disruptive source papers. On the other side, within “cites” edges, if the source (citation) and target (reference) paper have more identical citations, they may share more commonness and are more closely connected, leading to a positive effect. Unlike co-citing relationships, citations are expansive and continuously growing, making it challenging to differentiate disruptive papers based on co-cited relationships alone. Therefore, we employ a normalized score of co-cited strength that considers both similarity and accumulated count to identify highly-related developmental citations. It’s crucial to note that the co-cited/citing strengths are dynamic scores that evolve over time. As new papers emerge, the values of normalized co-cited/citing strengths undergo changes not only due to the direct influence of new citations but also because of alterations in numerical values and maximum values. Leveraging these co-cited/citing strengths can help uncover the dynamics inherent in multi-hop information diffusion. Overall, we introduce CompGAT, a novel module based on GATv2 (Brody et al., 2022), to encode comparative and co-cited/citing features between paper nodes. To capture the discrepancy between source and target nodes, we encode the concatenation of their representations. Besides the basic attention scores of GATv2, we incorporate co-cited or co-citing strengths by assessing the normalized similarity of their references or citations in the global citation network. Based on the adjacent matrix $\bm{A}$, it is calculated as $\bm{C}=\mathrm{norm}(\bm{A}\bm{A}^{T})$, which represents 2-hop relationships connected with intermediate paper nodes. Finally, the attention is a mixed distribution resembling CopyNet (Gu et al., 2016). For $i$-th paper node in $(l+1)$-th layer $\bm{h}_{i}^{(l+1)}$, we have:
@@
where $c_{ij}\in\bm{C}$ is the normalized co-cited/citing strength, $\lambda$ is to balance the sum distribution of co-cited/citing strengths and basic attention weights, and $\bm{w}_{a}^{(l)}$is the attention vector.
For other edges in the R-GCN, we apply GIN (Xu et al., 2019) with sum-pooling as the GNN encoder:
(5) $\begin{split}\bm{h}_{i,r_{dst}}^{(l+1)}=f_{\Theta,r}((1+\xi)\bm{h}_{i,r_{dst}}^{(l)}+AGG(\{\bm{h}_{j,r_{src}}^{(l)},&j\in\mathcal{N}(i,r_{src})\})),r\in\mathcal{R}^{normal}\end{split}$
where $\xi$ is the hyper-parameter to decide the retained original information, $\mathcal{R}^{normal}$ is the set including all edges in classic GIN, and $AGG$ is the neighborhood aggregation function of the GIN. In our model, we opt for sum-pooling as the aggregation method, as it can preserve crucial structural information including node degrees.
@@ 3.3.2. Type-specific Attention Snapshot Readout
To represent the time-aware information and filter noisy information of a single snapshot $\bm{h}_{t,s}$, we design an attention-based readout mechanism. It extracts the evolved temporal information of different types of paper nodes (reference, citation, and target paper) based on the publication age of the target paper.
(6) $\begin{split}&\bm{h}_{t,s}^{(l+1)}=\bm{h}_{t,s}^{(l)}+\sum_{j\in\mathcal{N}(t)}\hat{\alpha}_{tj}^{(l)}\bm{W}^{(l)}_{p}\bm{h}_{j,p}^{(l)}\\ &\hat{\alpha}_{tj}^{(l)}=\mathrm{softmax_{t}}(\hat{e}_{tj}^{(l)})\\ &\alpha_{tj}^{(l)}={\bm{w}_{a}}^{(l)}\mathrm{LeakyReLU}\left(\bm{W}^{(l)}_{s}\bm{h}_{t,s}^{(l)}+\bm{W}^{(l)}_{p}\bm{h}_{j,p}^{(l)}\right)+{\bm{w}_{t}}^{(l)}\mathrm{LeakyReLU}\left(f^{(l)}_{se}(\omega^{s}_{t})+f^{(l)}_{pe}(\omega^{p}_{j})\right)\end{split}$
where $\mathcal{N}(t)$ denotes the paper nodes within $t$-th 1-hop subgraph, $f^{(l)}_{se}(\omega^{s}_{t})$ and $f^{(l)}_{pe}(\omega^{p}_{j})$ are snapshot and paper type embeddings. The snapshot type is determined by binning the time interval between the time step $t$ and the publication time of the target paper. In this study, we categorize snapshot types based on temporal segments across all datasets: $\leq 5$ years, $5-10$ years, and $>10$ years. To enhance relevance across diverse fields of study, a potential refinement could involve tailoring these segments to align with the half-life of literature aging within specific domains.
3.3.3. Snapshot Information Mixer
To foster interactions between temporal and spatial dimensions, as well as between global and local information, we amalgamate the representations of paper nodes with the updated snapshot representation:
(7) $\begin{split}&\hat{\bm{h}}_{ti,paper}^{(l+1)}=mlp_{m}^{(l)}([\bm{h}_{ti,paper}^{(l+1)}\|\bm{h}_{t,s}^{(l+1)}])\end{split}$
where $mlp_{m}^{(l)}$ integrates and condenses the information from both the current paper node and the snapshot. leveraging the temporal information embedded in the snapshot representations through the mixer, we introduce extra temporal signals, enriching the structural encoder beyond sole reliance on graph structures.
@@ 3.3.4. Snapshot Temporal Encoder
To encode the temporal information across various snapshot times, we utilize a 4-layer Transformer Encoder. Compared with classic sequential models like RNN, the self-attention mechanism captures more intricate relationships that extend beyond strict ordinal dependencies:
@@ 3.3.5. Final Output
We obtain the target paper representation $\bm{o}$ by avg-pooling across the layers and then sum-pooling across the time steps, to maintain high-order temporal information:
@@ 3.4. Popularity-aware Disentanglement Module
Based on the encoded representation of the target paper from CGE, we design the Popularity-aware Disentanglement Module (PDM) to disentangle the citation count potentially received from different perspectives, including citation diffusion, collective conformity, and actual contribution. Overall, we sum up the individual predicted values to obtain the total citation increment and apply different auxiliary tasks to help the disentanglement. This approach enables the model to learn distinctive compositions for diverse papers, thereby facilitating the identification and utilization of their authentic contributions. We employ distinct Multilayer Perceptrons (MLPs) as encoders to process information within the target perspectives, and then feed them into another MLP separately to predict the corresponding citation increment $\hat{y}_{i}^{v}$:
@@
where $mlp^{d}_{enc}$ encodes specific aspects of the information and $mlp^{v}_{pred}$ predicts the citation increment from the target perspective. It is noteworthy that the three dedicated components are separately modeled to disentangle distinct perspectives, and they are trained simultaneously to facilitate comprehensive and concurrent learning.
@@ 3.4.1. Diffusion Encoder
The diffusion encoder extracts the popularity influence according to information diffusion, which is represented as diversity affected by spreading nodes’ degrees. We propose a Citation-aware Triplet Graph Contrastive Learning method to extract the trunk of spreading in the dynamic graph. In various snapshots, except for the target paper node, all other paper nodes are selectively omitted with probability, which is weighted by their normalized global citations within the local subgraph. Lowly-cited paper nodes are more likely to be dropped for the positive view $z^{pos}$, as they contribute less to spreading. Conversely, for the negative view $z^{neg}$, highly-cited paper nodes are more likely to be dropped, as they significantly impact the graph semantics. For every sample, we contrast them solely with their augmented views to preserve their unique characteristics from the diffusion perspective. We augment the subgraphs in every training epoch to enhance the robustness of learning. Through this approach, we amplify diffusion influence in the graph structure. Contrasted with original sample $z^{ori}$, the diffusion loss $\mathcal{L}_{dif}$ is:

### v4

@@ 3. Methodology
Figure 2. The overall architecture of DPPDCC. DPPDCC first encodes the Dynamic Heterogeneous Graph of the target paper with the Citation-aware GNN Encoder. Based on the encoded representation of the target paper, DPPDCC then disentangles its citation increment into diffusion, conformity, and contribution values through three corresponding auxiliary tasks.
In this section, we introduce our proposed model DPPDCC in detail. The overall framework of DPPDCC is depicted in Figure 2, which can be divided into two parts: Citation-aware GNN Encoder (CGE) and Popularity-aware Disentanglement Module (PDM).
@@ 3.1. Problem Formulation and Notations
In practice, the citation network comprises diverse entities and relations that evolve dynamically each year. To capture the entity-aware and time-variant characteristics, we construct a dynamic heterogeneous graph from the citation network for each target paper to predict its potential impact.
@@ 3.1.1. Dynamic Heterogeneous Graph
Given a target paper at the time step $t$, its heterogeneous subgraph is defined as $G^{t}=(\nu^{t},\epsilon^{t},\phi,\varphi)$. Here, $\nu^{t}$ denotes the set of nodes surrounding the target paper at the time step $t$. And $\epsilon^{t}$ represents the set of edges among $\nu^{t}$. Each node $\nu$ and edge $\epsilon$ are associated with their type mapping functions $\phi:\nu\rightarrow U$ and $\varphi:\epsilon\rightarrow R$, where $U$ and $R$ denote the types of nodes and edges shared across all times (see Figure 1c), respectively.
The backbone of the heterogeneous subgraph primarily consists of the cited and citing relations. We collect $k$-hop cited/citing neighbors of the target paper at the specific time $t$. For each node in $i$-th hop, to maintain the most recent and significant information, we select the $K_{i}$ newest published and then most highly cited papers as neighbors. To address the potential lack of global information, we incorporate metadata nodes associated with the papers (see Figure 1c). Prior to the observation time point, we consider a total of $T$ such subgraphs, each serving as a snapshot captured at distinct time steps. These subgraphs organized into a sequence collectively constitute the dynamic heterogeneous graph $G$ of the target paper.
@@ 3.1.2. Potential Impact Prediction
We estimate the potential impact of the target paper by predicting its citation increment, defined as the increase in citation count $\Delta$ years after the observation time point. By disentangling the sources of this increment, our model can more accurately assess the genuine potential impact, thereby providing a closer approximation to the real scientific significance.
@@ 3.2. Initialization
The initial embedding of the paper node is derived from the text embedding obtained by encoding the combination of title and abstract. Metadata node embeddings in each snapshot are the average embeddings over all papers associated with them in the global citation network at the corresponding time step. Specifically, the extraction of text embeddings is accomplished using Sentence-BERT (Reimers and Gurevych, 2019). To initialize the 00-th layer representation of the snapshot of the $i$-th target paper at the $t$-th time step $\bm{h}_{ti,s}^{(0)}$, we mainly take into account the contextual information from the corresponding time embedding $\bm{h}_{t,time}^{(0)}$ of time step $t$. Since the time embedding only preserves the global features, we utilize a gate module to approach specific information from the $i$-th target paper embedding $\bm{h}_{i,p}^{(0)}$. The snapshot representation $\bm{h}_{ti,s}^{(l)}$ of $l$-th layer will then be the proxy to aggregate the corresponding heterogeneous subgraph:
(1) $\begin{split}&\bm{h}_{ti,s}^{(0)}=\bm{a}_{ti}\times\bm{h}_{t,time}^{(0)}\\ &\bm{a}_{ti}=\mathrm{sigmoid}(\bm{W}^{input}\bm{h}_{t,time}^{(0)}+\bm{W}^{target}\bm{h}_{i,p}^{(0)})\end{split}$
@@ 3.3. Citation-aware GNN Encoder
Citation-aware GNN Encoder (CGE) aims to capture distinct structural and temporal features from the dynamic heterogeneous graph. To be specific, the backbone of CGE consists of modified R-GCNs (Schlichtkrull et al., 2018) as the structural encoder and a Transformer Encoder (Vaswani et al., 2017) as the temporal encoder. We stack $L$ integrated layers to encode them alternately instead of separate modeling. This allows us to exploit higher-order interactions between space and time. To elaborate, for subgraphs at different time points, we first utilize distinct Citation-aware R-GCNs to model their structural information respectively. Afterwards, the Type-specific Attention Snapshot Readout is employed to summarize the information from these subgraphs based on the snapshot representations. These aggregated representations are then fed into the temporal encoder to capture the intricate temporal relationships between different time points. By integrating structural and temporal encoding, our approach could extract valuable insights from the dynamic heterogeneous graph.
@@ 3.3.1. Citation-aware R-GCNs
Citation-aware R-GCNs are utilized to encode node hidden states $\bm{h}$ within the dynamic heterogeneous graph for structural modeling:
@@
where $\mathcal{R}$ is the set of all edges in the graph, $f_{r}$ is the message passing method applied to relation $r$, $g_{r}$ is the subgraph with relation $r$, $\bm{h}_{r_{src/dst}}$ is the hidden representation of the source/target node of relation $r$, and $AGG$ is the relation aggregation function.
Motivated by the Disruption-index (Wu et al., 2019), in ”is cited by” edges, if the source and target paper have more identical references, the source paper may be less influenced by the target paper, resulting in a negative effect. Conversely, in ”cites” edges, if the source and target paper have more identical citations, they may share more commonality and are more closely connected, leading to a positive effect. Thus, we introduce CompGAT, a novel module based on GATv2 (Brody et al., 2022), to encode comparative and co-cited/citing features between paper nodes. To capture the discrepancy between source and target nodes, we encode the concatenation of their representations. Besides the basic attention scores of GATv2, we incorporate co-cited or co-citing strengths by assessing the normalized similarity of their references or citations in the global citation network. Based on the adjacency matrix $\bm{A}$, it is calculated as $\bm{C}=\mathrm{norm}(\bm{A}\bm{A}^{T})$, which represents 2-hop relationships connected with intermediate paper nodes. Finally, the attention is a mixed distribution resembling CopyNet (Gu et al., 2016). For $i$-th paper node in $(l+1)$-th layer $\bm{h}_{i}^{(l+1)}$:
@@
where $c_{ij}\in\bm{C}$ is the normalized co-cited/citing strength, $\lambda$ is to balance the sum distribution of co-cited/citing strengths and basic attention weights, and $\bm{w}_{a}^{(l)}$is the attention vector. Additionally, we apply GIN (Xu et al., 2019) with sum-pooling for other edges in the R-GCN.
@@ 3.3.2. Type-specific Attention Snapshot Readout
To represent the time-aware information of a single snapshot $\bm{h}_{t,s}$, we design an attention-based readout mechanism. It extracts the evolved temporal information of different types of paper nodes (reference, citation, and target paper) based on the publication age of the target paper.
(4) $\begin{split}&\bm{h}_{t,s}^{(l+1)}=\bm{h}_{t,s}^{(l)}+\sum_{j\in\mathcal{N}(t)}\alpha_{tj}^{(l)}\bm{W}^{(l)}_{p}\bm{h}_{j,p}^{(l)}\\ &\alpha_{tj}^{(l)}=\mathrm{softmax_{t}}(\hat{e}_{tj}^{(l)})\\ &\hat{e}_{tj}^{(l)}={\bm{w}_{a}}^{(l)}\mathrm{LeakyReLU}\left(\bm{W}^{(l)}_{s}\bm{h}_{t,s}^{(l)}+\bm{W}^{(l)}_{p}\bm{h}_{j,p}^{(l)}\right)+e_{tj}^{(l)}\\ &e_{tj}^{(l)}={\bm{w}_{t}}^{(l)}\mathrm{LeakyReLU}\left(f^{(l)}_{se}(\omega^{s}_{t})+f^{(l)}_{pe}(\omega^{p}_{j})\right)\end{split}$
where $\mathcal{N}(t)$ denotes the paper nodes within $t$-th 1-hop subgraph, $f^{(l)}_{se}(\omega^{s}_{t})$ and $f^{(l)}_{pe}(\omega^{p}_{j})$ are snapshot and paper type embeddings. The snapshot type is determined by binning the time interval between time step $t$ and the publication time of the target paper.
@@ 3.3.3. Snapshot Temporal Encoder
To encode the temporal information within snapshot representations across various times $\bm{H}_{s}^{(l+1)}$, we utilize a multi-layer Transformer Encoder (Vaswani et al., 2017). Compared with classic sequential models, the self-attention mechanism captures more intricate relationships beyond strict ordinal dependencies:
@@ 3.3.4. Final Output
We obtain the target representation $\bm{o}$ by avg-pooling across the layers and then sum-pooling across the time steps to maintain high-order temporal information:
@@ 3.4. Popularity-aware Disentanglement Module
We design the Popularity-aware Disentanglement Module (PDM) to disentangle the citation count potentially received from different perspectives, including citation diffusion, collective conformity, and actual contribution. Overall, we train the three components simultaneously and sum up their predicted values to obtain the original increment. We further apply different auxiliary tasks to help the disentanglement. This approach enables the model to learn distinctive compositions for diverse papers, thereby facilitating the identification of their authentic contributions.
Specifically, we utilize separate Multilayer Perceptrons (MLPs) as encoders and predictors to handle information within the target perspectives. For the target perspective $v$, we begin by deriving its representation $\hat{\bm{o}}_{i}^{v}$ from the original target representation $\bm{o}_{i}$ using its encoder $mlp^{v}_{enc}$. Subsequently, we input this representation into its corresponding predictor $mlp^{v}_{pred}$ to anticipate its increment $\hat{y}_{i}^{v}$:
@@ 3.4.1. Diffusion Encoder
The diffusion encoder captures the influence of popularity based on information diffusion, which is characterized by diversity affected by the degrees of spreading nodes. To extract the trunk of spreading, we propose a Citation-aware Triplet Graph Contrastive Learning method. We employ citation/degree-aware node-dropping (Xu et al., 2022; He et al., 2023) as the graph augmentation technique. In different snapshots, apart from the target paper node, all other paper nodes are selectively omitted with probability, which is weighted by their normalized global citations within the local subgraph. Lowly-cited paper nodes are excluded for the positive view $z^{pos}$ since they contribute less to spreading. Conversely, for the negative view $z^{neg}$, highly-cited paper nodes are omitted, given their significant impact on the graph semantics. This approach amplifies the influence of diffusion in the graph structure. Contrasted with original sample $z^{ori}$, the diffusion loss $\mathcal{L}_{dif}$ is: